# Convergence of body-orders in linear atomic cluster expansions

### v1

@@ Abstract
We study the convergence of a linear atomic cluster expansion (ACE) potential [1] with respect to its basis functions, in terms of the effective two-body interactions of elemental Carbon systems. We build ACE potentials with descriptor sets truncated at body-orders $K=2,3,4$ trained on pure dimers, or on large datasets of different diversities but without any dimers. Potentials trained on a more diverse dataset fare better in validation and result in a nontrivial dimer curve, but still very far from the theoretical two-body interaction calculated by DFT. Moreover, dimer curves between descriptor sets clipped at different $K$ do not seem to converge to a universal function for a given dataset. We conclude that machine learning potentials employing atomic cluster expansions optimize losses at low $K$ but fail to generalize and converge properties described by two-body interactions.
@@ I Introduction
have been used as the framework to study material properties such as the thermodynamics of substitutional alloys [2, 3] and empirical potentials of condensed matter and molecular systems [4, 5, 6]. While the infinite expansion has been shown to be complete, its practical effectiveness is limited to cases where the total energy converges after a few body-orders $\nu$, as the number of terms increases combinatorially as higher orders are included.
@@
Developments on improving the construction of $E_{i}$ have been made in recent decades, with symmetry-respecting descriptors that use only a few terms in the expansion [7, 8, 9, 10] or take advantage of tensorial operations to reach an arbitrary body order [11, 1, 12, 13]. A detailed review on descriptors of local chemical environments has been done by Musil et al. [14]
While the completeness [2, 3, 15] of such expansions allow an apt description of the system, its form in Eq. 1 are considered to converge slowly [^1] Counterarguments [16] against the speed of convergence of body-ordered expansions claim that the surrounding environments are typically ignored and are usually crucial in the convergence at low body-orders. [4, 5, 17], where computing a term of order $\nu$ would scale in powers of $\nu-1$ of the number of neighbors. In order to avoid this prohibitive scaling, it is possible to express higher body-order terms as the product of simple two-body terms, at the price of including lower order terms, and losing the orthogonality of the expansion. A popular version of this resummation is the Atomic Cluster Expansion proposed by Drautz [1]. In this approach, each $E_{i}$ is expanded with a non-orthogonal basis, $\{A_{\nu}\}$, with unrestricted sums,
@@
and $\nu$ indexing the terms of the expansion.
The expansion has been shown to be complete [15], and implementations of the code have shown promise in machine learning applications [18, 19, 20, 21]. One of the state-of-the-art potentials, MACE [13], extends from the formalism of ACE and represents the state of the art in molecular dynamics applications for materials with multiple elements [22].
Whereas potentials adopting ACE descriptors have shown to converge fast in losses, a detailed study on the convergence of the body-orders with respect to properties described by lower orders in the original expansion in Eq. 1 have not been thoroughly investigated. In this work, we investigate the convergence of linear ACE at different body-orders $K$ in terms of their projection on the lowest terms of the original cluster expansion Eq. 1. In particular, we consider the convergence of the potential in energy and forces on different datasets, as a function of the number and type of basis functions utilized, and compare it to the stability of a real body ordered expansion. Our study focuses on the first nontrivial term of the expansion: the two-body term. This can be easily extracted from a given potential by studying the energy profile of two isolated atoms, a dimer, as a function of distance:
@@
Neural network potentials (NNPs) may also employ nonlinearities via functional embeddings [23, 24, 6], or are inherent from their activation layers[7, 25, 9, 10, 14]. While nonlinearities improve their performance by being able to represent higher-order terms when the expansion is truncated, one has to be careful in how they are applied as certain applications may invalidate the body-ordering of the descriptor [12]. We do not aim to build the most accurate potentials, but to have a more grounded understanding of their behavior in the ACE framework. As such, we limit our investigations to training simpler linear models as well as not going to too high of a body-order where returns are diminished. Linear ACE models perform comparably well relative to neural network potentials, and have been successfully employed to describe molecular [18] and condensed phases [19].
@@ II.1 Datasets
The datasets we use in the investigations consists of elemental Carbon structures. Three datasets were used: the first set consists of $247$ dimer structures at logarithmically-spaced distance between $0.5$ and $8$ Angstrom. Half of the structures were randomly selected for the validation set. The remaining two are sets of $1000$ structures coming from a collection of Carbon allotropes generated from [26]. One subset contains a homogeneous collection of diamond-like structures clustered using a Euclidean distance measure [27]. The diamond-like dataset is highlighted in the histogram of distances to a reference diamond structure in Fig. 1. The last subset are structures obtained using farthest point sampling [28] to maximize its diversity. For the diamond-like and diverse datasets, a validation set was constructed from $1000$ diverse structures that have not been seen in training. All structures studied energies and forces were calculated using the Quantum Espresso package [29] using the rVV10 functional [30] to account for nonlocal interactions.
Figure 1: Histogram of fingerprint distances of the configurations in the Carbon dataset in [26] relative to a reference diamond-like structure (inset). $1000$ configurations were chosen within a narrow region in structure space.
@@ II.2 Descriptor
We use the ACE descriptor as implemented by the pacemaker package [20]. We used scaled Chebyshev polynomials as the radial functions, vanishing smoothly at the cutoff length of $r_{\text{cut}}=5$. We start from $K=2$ body functions and progressively add more terms of the same order. We then introduce higher order $K=3$ terms while keeping the number of lower order terms fixed, and then repeat the procedure for higher orders. ACE descriptors employ spherical harmonics to describe angular functions, with $n,l$ values chosen to respect maintain rotational symmetry. Low $n,l$ values were chosen to avoid overfitting the model from too many basis functions. We note that the original ACE and the current version of pacemaker include self-interactions from $K>2$ body orders, so redundancies of lower-body order functions appear as functions of higher $K$.
@@ II.3 Fitting procedure
We use the default loss function $\mathcal{L}$ of pacemaker [20],
$\mathcal{L}=\left(1-\kappa\right)\Delta^{2}_{E}+\kappa\Delta^{2}_{F}+\Delta_{\text{coeff}}+\Delta_{\text{rad}}$ (6)
where $\Delta^{2}_{E}$ and $\Delta^{2}_{F}$ are the corresponding energy and force root-mean-square errors (RMSE), and $\kappa$ is the coefficient that balances the energy and force contributions to $\mathcal{L}$. The term $\Delta_{\text{coeff}}$ regularizes the basis function coefficients [^2] This term includes the L1 and L2 regularizations., while $\Delta_{\text{rad}}$ regularizes the coefficients of the radial components of the basis function. The dimer dataset was randomly split between training and validation set, while the diamond-like and generic datasets were validated using a set similar to the generic set. To reconstruct the 2-body contribution to the cluster expansion, we build the energy curve of two isolated Carbon atoms as in Eq. 5. The converged potentials from pacemaker were evaluated using the Atomic Structure Environment (ASE) code [31].
@@ III Results and Discussion
Figure 2: Top: $\mathcal{L}$ as a function of number of basis functions, $N_{f}$, for the dimer dataset. Training and validation errors as dashed and solid curves. Curve colors indicate potentials built from basis sets which are exclusively two-body (light green), up to three-body (green), and up to four-body functions (dark green), respectively. Light solid lines indicate $\mathcal{L}$ if more functions introduced keeping the previous $K=2$ or $K=3$. Bottom: dimer curves for each of the potentials trained at certain $N_{f}$, with reference DFT data as points. Heavy lines represent the potential with the largest $N_{f}$ before functions of higher $K$ are introduced.
We start by fitting the dimer dataset with an increasing number of basis functions, $N_{f}$, up to $250$ elements and $K=4$ body-order. The loss and dimer curves are shown in Fig. 2, where light green, green, and dark green indicate data from potentials trained using $K$ up to 2, 3, and 4, respectively. The top plot shows the dependence of the loss $\mathcal{L}$ from Eq. 6 to $N_{f}$. The validation loss (solid curve) converges $K=2$ and increases with the addition of higher order functions, jumping two orders of magnitude as more $K=4$ functions are added. This is of course to be expected, as the dataset consists solely of 2 body structures. The bottom plot highlights the dimer curves of the potentials trained from the dimer set at the highest $N_{f}$ of a particular order $K$. We see that all potentials were able to reproduce the dimer curve as expected from the DFT results.
The energy RMSE for the representative potentials were around $5$ meV/at, and force RMSE were slightly less than $200$ meV/Å for $K=2$ and $K=3$. The force RMSEs increased to $2700$ meV/Å for the overfitting potential that at $N_{f}=250$, and are due to the short-range region of the curve having energies up to almost two orders of magnitude higher, making the deviations from DFT results more sensitive to errors. A comparison of the forces at the short range region may be found in the Supporting Information B.1.
@@
Figure 3: $\mathcal{L}$ vs. $N_{f}$ (top) and dimer curves (bottom) for the homogeneous, diamond-like dataset. Orange and pink plots correspond to potentials with only $K=2$ and up to $K=3$ functions, respectively while purple curves include $K=4$ functions. The x-values of the dimer curve start at the minimum distance found in the homogeneous dataset.
@@
We then look at the dimer curves produced by the diverse dataset for different descriptor complexities, ranging from $K=2$ to $5$, with various sizes for $K=4$ (similar loss analysis with consistent earlier dimer curves are shown in the Supporting Information B.2.). Fig. 4 shows these curves having more pronounced shapes as compared to the ones produced by the diamond-like dataset. While the curves look similar as more functions of the same $K$ are introduced in the set as seen in the $K=4$ potentials, the curves of different $K$ limits do not seem to converge even when five-body contributions are already added. The variability is much more pronounced, even among potentials with similar maximum $K$ and more so when adding new terms. We posit that the effective two-body potential does not seem to converge, even for when five-body basis functions are added, leading to an arguably well-converged potential.
, Refer to caption
Figure 4: Dimer curves for the diverse $1000$-structure dataset for extended descriptor sizes. While curves produced by models within a certain $K$ ($K=4$) may appear to converge, they qualitatively do not seem to approach a consistent shape whenever functions of higher $K$ are introduced in the descriptor set.
Lastly, we train a potential that includes the $50000$ structures on the whole Carbon dataset in [26] to see if further increasing the complexity of the input data affects our conclusion. The dimer curves shown in Fig. 5 tell the same story as the one of smaller dataset: the curves do not approach the DFT dimer results at all, and they do not seem to converge to a fixed function as higher $K$ basis elements are introduced.
It is worth reminding that we are testing the generalization capabilites of the diverse set potentials and not their ability to capture in-domain data. The pair-interactions present in the set are those that exist in surrounding materials and not in isolation. A brief test shown in Supporting Information B.3. shows that gradually populating the diverse dataset with dimers eventually leads to a binding curve that approaches the DFT results.
Figure 5: Dimer curves for expanded $50000$-structure dataset for multiple descriptor sizes. Like the potentials trained with a smaller dataset in Fig. 4, potentials trained with increasing descriptor sizes but within a certain $K$ seem to converge, but introducing the descriptor set with $K=5$ basis functions vary the shape considerably.
@@
Table 1: Comparison of linear ACE models to neural network potentials in MAE for energies and forces for the Carbon dataset in [26] for datasets containing $n_{\text{set}}=1000$ and $50000$ structures. Energies (E) are in meV/atom and force components (F) are in meV/Å.
@@
Potentials E F E F
$K=4,250N_{f}$ 48.50 241.27 44.06 222.42
$K=4,493N_{f}$ 18.30 199.23 17.55 176.48
$K=4,930N_{f}$ 15.40 167.73 15.27 166.70
$K=5,793N_{f}$ 14.88 160.78 14.74 159.80
@@
We also make sure that the performance of our linear models are reasonable by comparing their loss metrics to the performance of different neural network potentials trained on the same dataset in [32] in Table 1. For the $1000$-size subsets, we found that the linear ACE models with the most complex descriptor sets ($K=4,N_{f}=930$ and $K=5,N_{f}=793$) were able to outperform all PANNA potentials and even the $l=1$ version of the NequIP [33] potential. However, they perform relatively worse with all other potentials when the complete $50000$-size Carbon set is used as the training set, which suggests that the nonlinearities absent in the ACE models provide a considerable improvement in training more complex datasets. Despite this, we assert that the inability of the linear models to capture the effective two-body interaction and converge to a defined shape are not due to the poor $\mathcal{L}$, but rather the nature and truncation of the expansion employed in the descriptor sets.

### v2

@@ Abstract
We study the convergence of a linear atomic cluster expansion (ACE) potential [1] with respect to its basis functions, in terms of the effective two-body interactions of elemental Carbon and Silicon systems. We build ACE potentials with descriptor sets truncated at body-orders $K=2$ to $K=5$ trained on a diverse Carbon dataset and on Silicon dimers to pentamers. The potentials trained on diverse structures with standard ACE bases are not able to recover the correct dimer curves much less produce stable curves or solutions. While employing ACE bases removed of self-interactions still does not generalize to the DFT-expected results, properly tailored datasets and basis sets are able to show signs of convergence and stability in the curves and expansions, suggesting a method to build potentials with interpretable bases with respect to the cluster expansion.
@@ I Introduction
have been used as the framework to study material properties such as the thermodynamics of substitutional alloys [2, 3] and empirical potentials of condensed matter and molecular systems [4, 5, 6]. While the expansion has been shown to be complete [2, 3, 7] when the number of atoms matches the highest body order of the expansion, its practical effectiveness is limited to cases where the total energy converges after a few body-orders $\nu$[4, 5, 8], as the number of terms increases combinatorially as higher orders are included.
@@
Improvements on the construction of $E_{i}$ have been made in recent decades, with symmetry-respecting descriptors that use only a few terms in the expansion [9, 10, 11, 12] or take advantage of tensorial operations to reach an arbitrary body order [13, 1, 14, 15]. A detailed review on descriptors of local chemical environments has been done by Musil et al. [16]
A popular version of this resummation is the Atomic Cluster Expansion proposed by Drautz [1]. In this approach, each $E_{i}$ is expanded with a non-orthogonal basis, $\{A_{\nu}\}$, with unrestricted sums,
@@
and $\nu$ indexing the terms of the expansion. The expansion has been shown to be complete [7], and its implementations are effective in machine learning applications [17, 18, 19, 20, 21]. One of the state-of-the-art potentials, MACE [15], extends the formalism of ACE as a graph neural network but the interpretability of the expansion is traded for better expressivity when compared to linear potentials.
In employing interatomic potentials using the ACE framework, the selection of the basis complexity in the number of functions and maximum body-order $K$ is a heuristic process and no systematic construction has been proposed yet. While ACE potentials have been shown to achieve low fitting errors, the stability of these solutions and the interpretability of the basis functions are not as studied, beyond undertakings to eliminate the self-interactions in Ho et al. [22] and perform uncertainty quantifications in linear models in Chong et al. [23].
In the advent of foundational machine learning models [24] as baselines training neural network potentials that do not need to be trained from scratch, it may also be worth asking if a method to systematically build a body-ordered basis with reasonable lower-order energetics can be established. In this work, we investigate the convergence of linear ACE at different body-orders $K$ in terms of their projection on the lowest terms of the original cluster expansion Eq. 1. In particular, we consider the convergence of the potential in energy and forces on different datasets as a function of the number and type of basis functions utilized, and compare it to the stability of a real body-ordered expansion. Our study focuses on the first nontrivial term of the expansion: the two-body term. This can be easily extracted from a given potential by studying the energy profile of two isolated atoms, a dimer, as a function of distance:
@@
While nonlinearities can improve the performance of machine learning potentials by representing higher-order terms when the expansion is truncated, care should be taken in how they are applied as certain applications may invalidate the body-ordering of the descriptor [14]. In the present work, we do not aim to build the most accurate potentials, but to have a more grounded understanding of their behavior in the ACE framework. As such, we limit our investigations to training simpler linear models as well as not employing very high of a body-order where returns are diminished. Linear ACE models perform comparably well relative to neural network potentials, and have been successfully employed to describe molecular [17] and condensed phases [18].
@@ II.1 Datasets
The datasets we use in the investigations consists of elemental structures of Carbon and Silicon. For Carbon, three datasets were used: the first set consists of 247247247 dimer structures at distances between 0.50.50.5 Å and 888 Å, with half selected for the validation set. The other two are sets of 100010001000 structures coming from a collection of Carbon allotropes at low pressures (00.10-0.10 - 0.1 GPa) [25]. One subset contains a homogeneous collection of diamond-like structures clustered using a Euclidean distance measure [26]. The diamond-like dataset is characterized through a histogram of distances to a reference diamond structure in Fig. 1. The last subset are structures obtained using farthest point sampling [27] to maximize its diversity. For the diamond-like and diverse datasets, a validation set was constructed from 100010001000 diverse structures that have not been seen in training. The Carbon data were calculated using the Quantum Espresso package [28] using the rVV10 functional [29] to account for nonlocal interactions.
Figure 1: Histogram of fingerprint distances of the configurations in the Carbon dataset in [25] relative to a reference diamond-like structure (inset). 100010001000 configurations were chosen within a narrow region in structure space.
[…]
@@ II.2 Descriptor
We use a linear ACE descriptor as implemented by both the pacemaker package [19] and the ACEpotentials.jl [21]. We start from $K=2$ body functions and progressively add more terms of the same order. We validate the models at each addition of basis functions, and monitor when the interim model begins to overfit. Before the generalization error increases, we then introduce higher order $K=3$ terms while keeping the number of lower order terms fixed, and then repeat the procedure for higher orders. ACE descriptors employ spherical harmonics to describe angular functions, with $n,l$ values chosen to respect rotational symmetry. Low $n,l$ values were chosen to avoid overfitting the model from too many basis functions.
@@ II.3 Fitting procedure
Training was done with a least squares error method for both pacemaker and ACEpotentials.jl methods. For pacemaker the loss $\mathcal{L}_{p}$ is,
$\mathcal{L}_{p}=\left(1-\kappa\right)\Delta^{2}_{E}+\kappa\Delta^{2}_{F}+\Delta_{\text{coeff}}+\Delta_{\text{rad}}$ (6)
where $\Delta^{2}_{E}$ and $\Delta^{2}_{F}$ are the corresponding energy and force root-mean-square errors (RMSE), and $\kappa$ is the coefficient that balances the energy and force contributions to $\mathcal{L}$. The term $\Delta_{\text{coeff}}$ regularizes the basis function coefficients [^1] This term includes the L1 and L2 regularizations., while $\Delta_{\text{rad}}$ regularizes the coefficients of the radial components of the basis function. ACEpotentials.jl minimizes the loss $\mathcal{L}_{A}$ through the kernel ridge regression methods,
$\mathcal{L}_{A}=||\mathbf{W}\left(\mathbf{y}-\mathbf{Ac}\right)||^{2}+\lambda||\mathbf{\Gamma c}||^{2}$ (7)
where $\mathbf{W}$ encodes the relative loss weightings for energies and forces, $\mathbf{y}$ contains the labeled data, $\mathbf{A}$ is the basis matrix and $\mathbf{c}$ are the parameters to be learned to minimize $\mathcal{L}_{A}$. $\mathbf{\Gamma}$ contains a more general treatment to regularization allowing different penalizing strengths to the coefficients $c$, and $\lambda$ dictates the strength of the regularization. We used the Bayesian Linear Regression to let the model dictate appropriate values for $\lambda$ and $\mathbf{\Gamma}$. Details of the regression methods are found in Witt et al. [21], and the Supporting Information discusses more details in the numerical setup. Evaluation of dimer curves were done using the Atomic Simulation Environment (ASE) by Hjorth Larsen et al. [30].
Aside from the parameters involving basis set complexities, radial bases/geometric priors, and the loss weightings, other values were set to default for both pacemkaer and ACEpotentials.jl to ensure a reasonable set of controlled parameters.
@@ III Results and Discussion
III.1 Self-interacting ACE bases
[…]
Figure 2: pacemaker— (Top) Loss as a function of number of basis functions, $N_{f}$, for the dimer dataset. Training and validation errors as dashed and solid curves. Curve colors indicate potentials built from basis sets which are exclusively two-body (light green), up to three-body (green), and up to four-body functions (dark green), respectively. Light solid lines indicate the loss if more functions were introduced without introducing higher $K$. (Bottom) dimer curves for each of the potentials trained at certain $N_{f}$, with reference DFT data as points. Heavy lines represent the potential with the largest $N_{f}$ before functions of higher $K$ are introduced.
We start by fitting the dimer dataset with an increasing number of basis functions, $N_{f}$, up to 250250250 elements and $K=4$ body-order. The loss and dimer curves are shown in Fig. 2, where light green, green, and dark green indicate data from potentials trained using $K$ up to 2, 3, and 4, respectively. The top plot shows the dependence of the loss function $\mathcal{L}$ on $N_{f}$. The validation loss (solid curve) converges for $K=2$ and increases with the addition of higher order functions, jumping two orders of magnitude as more $K=4$ functions are added. This is of course to be expected, as the dataset consists solely of 2 body structures. The bottom plot highlights the dimer curves of the potentials trained from the dimer set at the highest $N_{f}$ of a particular order $K$. We see that all potentials are able to reproduce the dimer curve as expected from the DFT results.
The energy RMSE for the representative potentials are around 555 meV/at, and force RMSE are slightly less than 200200200 meV/Å for $K=2$ and $K=3$. The force RMSEs increase to 270027002700 meV/Å for the overfitting potential at $N_{f}=250$, and are due to the short-range region of the curve having energies up to almost two orders of magnitude higher, making the deviations from DFT results more sensitive to errors. A comparison of the forces at the short range region may be found in the Supporting Information B.1.
@@ III.1 Self-interacting ACE bases
Figure 3: pacemaker— (Top) $\mathcal{L}$ vs. $N_{f}$ and (Bottom) dimer curves for the homogeneous, diamond-like dataset. Orange and pink plots correspond to potentials with only $K=2$ and up to $K=3$ functions, respectively while purple curves include $K=4$ functions. The x-values of the dimer curve start at the minimum distance found in the homogeneous dataset.
@@
We then look at the dimer curves of the potentials for different descriptor complexities, ranging from $K=2$ to 555. The blue-green curves are potentials trained with 100010001000-structures while the orange-red curves are trained with a larger 500005000050000-structure set. Fig. 4 shows these curves having higher variability compared to the ones produced by the diamond-like dataset. While the curves look similar as more functions of the same $K$ are introduced in the set, as seen in the $K=4$ potentials, the curves of different $K$ limits do not seem to converge even when five-body contributions are already added. The variability is much more pronounced, even among potentials with similar maximum $K$ and more so when adding new terms.
[…]
Figure 4: pacemaker— Dimer curves for the diverse 100010001000-structure (blue-green) and 500005000050000-structure (orange-red) for various descriptor sizes. No convergence observed for both sets as higher $K$ functions are introduced.
We note that the lack of pairs around the 2.02.02.0 Å region are due to the structures in the dataset belonging to relatively low pressures. The comparisons between the 100010001000-structure and the more complete 500005000050000-structure datasets have virtually similar pair distribution functions due to the farthest point sampling methods done to sample the smaller subset. The pair distribution comparisons may also be seen in Supporting Information C.
Finally, we try the same experiment using ACEpotentials.jl using the 100010001000-structure set with the plots. Fig. 5 shows qualitative improvements of the dimer curve in the equilibrium bond length $r_{0}$ and energy minimum may be attributed to ACEpotentials.jl requiring additional user inputs (i.e. geometric priors) and more rigorous regularization schemes[21]. The routines involving Bayesian Linear Regression also allow for the automatic selection of regularization strengths $\lambda$ in Eq. 7 which helps reduce the number of hyperparameter sweeps done. In contrast, the absence of a more involved radial basis construction [^2] The radial basis construction of pacemaker focuses on inner cutoffs that penalize oscillatory behavior at the short-range regions where data may be lacking, but optimizing this region of the curve is outside the study scope. in the qualitative features of the dimer curve lead to its greater reliance on a richer dataset.
Still, it seems that the potentials made from purified bases (solid curves) do not produce a marked improvement over the self-interacting basis (dashed lines) as they are mostly overlapping. While it seems that no improvements are seen in the purified basis, the study of the gram matrix conditioning from [22] indicates that the potentials become less stable when the total number of relevant particles $J$ (or neighbors in periodic systems) far exceed the maximum body-order, $K$, of the basis used to describe the potential [^3] The notation of Ho et al. uses $N$ for the body-order variable instead of $K$, and uses $J$ to indicate the number of particles or neighbors considered in the system..
[…]
Figure 5: ACEpotentials.jl— Dimer curves for 100010001000-structure Carbon dataset trained with the Julia code [21]. Dashed lines indicate the ACE basis ($K=3,4,5$) with self-interactions similar to the one in pacemaker, while the solid lines are trained with a basis purified of self-interactions. Gray bars show the pair distance distribution in the training set. Using a purified basis for the Carbon dataset yields practically identical dimer curves with those from the self-interacting basis for up to $K=4$, with a qualtitative difference in the energy well for the $K=5$ potential.
@@
Table 1: Comparison of linear ACE models to neural network potentials in MAE for energies and forces for the Carbon dataset in [25] for datasets containing $n_{\text{set}}=1000$ and 500005000050000 structures. Energies (E) are in meV/atom and force components (F) are in meV/Å.
@@
Linear Potentials E F E F
pacemaker
$K=4,N_{f}=250$ 48.50 241.27 44.06 222.42
$K=4,N_{f}=493$ 18.30 199.23 17.55 176.48
$K=4,N_{f}=930$ 15.40 167.73 15.27 166.70
$K=5,N_{f}=793$ 14.88 160.78 14.74 159.80
ACEpotentials.jl
$K=3,N_{f}=297$ 38.10 336.84 - -
$K=4,N_{f}=802$ 18.00 210.14 - -
$K=5,N_{f}=929$ 16.72 202.78 - -
Neural Networks
@@
We also make sure that the performance of our linear models are reasonable by comparing their loss metrics to the performance of different neural network potentials trained on the same dataset in Pellegrini et al. [31], as shown in Table 1. For the 100010001000-size subsets, we found that the linear ACE models with the most complex descriptor sets ($K=4,N_{f}=930$ and $K=5,N_{f}=793$) were able to outperform all PANNA potentials and even the $l=1$ version of the NequIP [32] potential. While promising, we may attribute the poor performance of the neural network potentials from the fact that 100010001000 structures may not be enough data and the resulting potentials may overfit without appropriate regularizations.
However, they perform relatively worse with all other potentials when the complete 500005000050000-size Carbon set is used as the training set, which suggests that the nonlinearities absent in the ACE models provide a considerable improvement in training more complex datasets. Despite this, we believe that the inability of the linear models to capture the effective two-body interaction and converge to a defined shape are not due to the poor $\mathcal{L}$, but rather the nature and truncation of the expansion employed in the descriptor sets, which is further discussed in the experiments with purified bases for the Silicon dataset.
III.2 Purified ACE bases
We then repeat the same investigation on a dataset of Silicon structures[23] containing dimers, trimers, tetramers, and pentamers so the maximum number of atomic neighbors matches $K=5$. We train potentials from subsets of 100100100 trimers (Si3), 100100100 tetramers (Si4), and 500500500 pentamers (Si5), and a mixture of the structures leaving out the dimers (Si345) to mimic the diverse Carbon set. Fig. 6 highlights that the out-of-distribution Silicon potentials, those trained without dimers, also have curves that are not qualitatively recovered. Additionally, the Si345 potentials, which strikes the closest parallel to the diverse Carbon dataset, exhibit the overlapping of purified and self-interacting potentials albeit at a lesser degree than the Carbon tests. When we look at the coefficients of the expansion, the non-overlap of the lines show the instability of the solutions for the potentials at different $K$.
When focusing on the purely trimer trained Si3 potentials, we see signatures of convergence at the purified basis. Dimer curves of potentials at all $K$ overlapping and the basis coefficients $c_{pure}$ for the 222- and 333- body functions (in the blue region) are consistent with each other. The $c_{pure}$ corresponding to $K\geq 4$ functions are 0, which suggest that the self-interactions are indeed removed. In the Si4-trained potentials, dimer curves for the $K=3$ basis set already deviates from the others due to the lower-order $c_{pure}$ misrepresenting the higher-order tetramer interactions. We also see this happening for the Si5-trained potentials and even for the $K=4$ basis sets for their misrepresentation of the 555-body interactions. Overall, the qualitative features of the dimer curves do not seem to converge as basis functions become more complex with higher $K$. This makes it difficult to apply the potentials to structures it has not seen during training.
[…]
Figure 6: ACEpotentials.jl— (Bottom of each quadrant) Dimer curves of potentials from the Silicon structures in Chong et al. [23] trained at different subsets. Simn indicate potentials were trained from datasets that include $m$-mers and $n$ mers, black points show the DFT dimer energies, and gray bars indicate the distributions of pairs in the dataset. Overlapping solid orange ($K=4$) and green ($K=5$) lines indicate that potentials with purified bases eventually converge to a defined shape when the total number of particles $N$ match the maximum body order $K$. (Upper plots of each quadrant) Coefficients of the purified ($c_{pure}$) and self-interacting ($c_{self}$) potenials. Blue, orange, and green lines correspond to coefficients of potentials capped at $K=3$, $K=4$, and $K=5$ body-ordered functions while blue, orange, and green shaded regions indicate coefficients corresponding to $K=2,3$, $K=4$, and $K=5$ body-ordered functions. Overlapping lines for a certain region mean that potentials have very similar basis coefficients, suggesting the expansion has stabilized for that particular body-order.
While generalization to the correct DFT curve seems unlikely, we can still see how the coefficients of purified linear ACE models perform when doing interpolative tasks and dimers are included in the training set. It is evident in Fig. 7 that the correct DFT results are recovered regardless of basis max $K$ or whether the basis was purified or not. To understand whether the solution we obtain is stable, we again look at the coefficients of the potentials and see that the purification is successful in disentangling lower-order terms from the contributions of functions at higher $K$. It is also worth noting that the $c_{self}$ for the Si2-trained datasets look identical to the purified coefficients, and is possibly due to the fact that the self-interaction terms are not needed to capture the dimer energetics.
[…]
Figure 7: ACEpotentials.jl— (Top of each quadrant) ACE coefficients and (bottom of each quadrant) dimer curves of Silicon potentials trained at different subsets, including dimers in all of them. All dimer curves recover the correct DFT values, and purified potentials. All $c_{self}$, except for the Si2 training, had higher-order renormalizations as seen in the nonzero $K=3,4$ for Si23 and $K=4$ for Si234. $c_{pure}$ coefficients are similar for the $K=2,3$ terms in Si23-trained potentials, and similar for $K=4$ terms in the Si234 potentials.
The mismatch in the overlap of $c_{pure}$ can still be seen in the $K=2,3$ functions of the Si234 sets and the $K=2,3,4$ functions of the Si2345 sets which is the signature of aliasing that occurs when the basis is not complex enough to capture the higher-order energetics. To see if we are consistently getting not only the correct dimer curve, but a stable pair potential, we look at the $c_{pure}$ of the $K=2$ basis functions for datasets where max $K$ matches the maximum number of particles $J$ in Fig 8. We include the Si2 potential with the $K=3$ basis as the reference coefficient set. Surprisingly, we see that no other potential was able to capture the reference coefficients completely, with greater deviations coming from the Si234 and Si2345 potentials. Looking at the discrepancy between training and validation errors between potentials, the high validation error for the Si234 potential suggests that more tetramers need to be sampled. For the Si2345 potential, both training and validation errors being high implies that the $K=5$ basis functions as well as more pentamers are needed.
[…]
Figure 8: ACEpotentials.jl— $c_{pure}$ of the $K=2$ basis functions for potentials with basis complexities matching the max number of particles, $K=J$.